### v1

# Duoduo CLIP: Efficient 3D Understanding with Multi-View Images

Abstract
We introduce Duoduo CLIP, a model for 3D representation learning that learns shape encodings from multi-view images instead of point-clouds. The choice of multi-view images allows us to leverage 2D priors from off-the-shelf CLIP models to facilitate fine-tuning with 3D data. Our approach not only shows better generalization compared to existing point cloud methods, but also reduces GPU requirements and training time. In addition, we modify the model with cross-view attention to leverage information across multiple frames of the object which further boosts performance. Compared to the current SOTA point cloud method that requires 480 A100 hours to train 1 billion model parameters we only require 57 A5000 hours and 87 million parameters. Multi-view images also provide more flexibility in use cases compared to point clouds. This includes being able to encode objects with a variable number of images, with better performance when more views are used. This is in contrast to point cloud based methods, where an entire scan or model of an object is required. We showcase this flexibility with object retrieval from images of real-world objects. Our model also achieves better performance in more fine-grained text to shape retrieval, demonstrating better text-and-shape alignment than point cloud based models.
1 Introduction
Learning representations for 3D shapes is increasingly useful with growing applications in object generation, shape understanding, grounding and robotics. In recent years, there has also been growing interest in 3D representations that are semantically aligned with text. By having a 3D representation aligned to language, these representations can easily be used for text-to-3D shape retrieval [36, 41, 45] and generation [4], or as a basis for 3D feature maps that can be queried with text [16, 25, 19].
Initial work in this domain [4, 36, 41] attempted to train such aligned models on limited data. More recent work [49, 25] leveraged large pre-trained vision-language models (e.g. CLIP [35]) that already align the text and image embeddings. ULIP [49] used frozen CLIP text-image encoders, and trained only the point-cloud encoder but was still limited due to the limited number of 3D objects as well as paired text captions. With larger 3D datasets like Objaverse [7] becoming available in conjunction with automated captioning methods like BLIP-2 [23], recent works [50, 25, 33, 60] trained on such datasets have showcased better performance and stronger generalizability. Despite this, more and more GPU resources are required for training such methods. These recent works [50, 25, 33] also predominantly choose point clouds as their 3D representation. The point clouds are typically limited in resolution and have a domain gap with real-world images, and thus these point cloud based methods can not effectively leverage the 2D priors from CLIP. Uni3D [60] proposed a training regime for incorporating 2D priors with point cloud as input. However, their model requires significant scaling (1B parameters) for the best performance. In this work, we instead choose multi-view images as our shape representation. This not only allows for an easier extension and a more natural way to adapt off-the-shelf 2D representation models. Our Duoduo CLIP also results in more efficient training. As shown in Figure 1(b), our model not only has much stronger generalization on shapes that were not seen during training with a modest model size, but also significantly lowers the amount of compute resources required for training (see Table 1). While some recent work [39, 24] also explore using multi-images with CLIP, their methods do not investigate the use of contrastive learning and only focus on zero/few-shot performance on small-scale datasets.
(a) We use multi-view images, instead of point-clouds, to represent the shape.
(b) Accuracy on Objaverse-LVIS with all models trained on Ensembled (no LVIS). Our Duoduo CLIP is fast to train, relatively small while also achieves the best performance.
Figure 1: We propose Duoduo CLIP, a simple model for 3D shape understanding that represents the shape with a multi-view images encoder instead of a point-cloud (PC) encoder. Compared to recent PC-based methods (e.g. OpenShape [25], Uni3D [60]), our model is both accurate and efficient.
Table 1: Comparison of training time and GPU used for our method and recent point cloud based methods.
Method GPU Time
OpenShape [25] 1$\times$×A100 (80GB) 300 hr
Uni3D [60] 24$\times$×A100 (40GB) 20 hr
RECON++ [33] 8$\times$×A800 (80GB) 1 day
Ours (Full) 4$\times$×A40 (48GB) 14.3 hr
Ours (6 layers) 4$\times$×A5000 (24GB) 14.3 hr
Using multi-view images also offers greater flexibility in many scenarios. Point cloud based methods require a fairly complete scan of the object, which can be challenging to obtain. In contrast, our model can encode objects from a variable number of frames (1 to M), with more frames capturing more details. Note that we use frames and views interchangeably. This makes our model plug-and-play for any method using image features for encoding 3D objects, enhancing performance with multiple frames. Additionally, previous methods that only train on synthetic objects have limited performance on real-world objects. We further train both our model point cloud based methods on real photos and densely reconstructed objects. Due to artifacts in the point cloud reconstruction pipeline, our model outperforms on real-world objects. Performance is boosted with additional multi-view images where point clouds aren’t available, demonstrating our model’s potential for scaling as more data becomes available.
In summary, our Duoduo CLIP model provides the following: Simple model. We average embeddings across views of an object to obtain 3D representations, while modifying the attention layers to span multiple frames of the shape to enable information flow across frames in the model. Better efficiency. We explore several strategies for efficiently fine-tuning off-the-shelf CLIP models, including training only the attention layers and selectively choosing the number of layers to train without hurting performance. More flexible representation. We showcase benefits of choosing multi-view images as a more flexible 3D representation over point clouds on real-world objects. Furthermore, we also demonstrate better text to shape retrieval with more fine-grained descriptions.
2 Related Work
There is a large literature of work on 3D representation learning from 3D data either using supervised training [31, 34] or using self-supervised learning (with reconstruction loss [5], contrastive learning [48], masked prediction [54], and others). Here we focus on work that learns aligned representation spaces between text and either 2D or 3D.
2D-text joint representation. Modeling multiple modalities through a shared representation has been explored through early works [10, 11, 21, 47] by aligning images and text embeddings. With the introduction of the InfoNCE loss [29], ConVIRT [57] proposed using contrastive learning with InfoNCE style loss to learn joint text and image embeddings. This was popularized by follow-up work such as CLIP [35] and ALIGN [20], which demonstrated that by scaling up training on a large-scale dataset, the learned embeddings can be useful for many downstream tasks such as classification, retrieval, grounding and generation. Followup work [20, 51, 28, 55, 52] has since been proposed to improve upon the formulation or training. There have also been attempts such as MERU [8] that use a hyperbolic embedding space to better model the hierarchical relations between text and images.
3D-text joint representation. Text2Shape [4] was one of the first works that introduced a captioned text and 3D dataset and used metric learning to learn an aligned space between text and 3D shapes. Follow-up work investigated how to improve the aligned representation for text-to-shape retrieval with triplet loss [41] and contrastive losses [36, 49, 45]. Prior work also investigated the use of CLIP to align the embeddings [43, 36, 49]. ULIP [49] used frozen CLIP text-image encoders, and trained only the point-cloud encoder. However, the generalizability of these works were limited by the available dataset size at the time. With the introduction of larger scale 3D datasets such as Objaverse [7], and the use of LVLMs for captioning 3D datasets [25, 27], more recent work [25, 50] were able to train with larger datasets of paired text and 3D shapes. OpenShape [25] and ULIP-2 [50] were among the first to use such datasets to train a point cloud encoder aligned with text using contrastive learning. TAMM [58] leverages additional adapters to better align with the pre-trained CLIP models. Uni3D [60] exploits 2D priors within pre-trained large-scale vision models and maps point cloud patches to image patches within pre-trained ViTs to enable the scaling of even larger models. VIT-LENS [22] follows a similar pipeline by training a point cloud embedding layer and perceiver model to map to a frozen CLIP model. RECON++ [33] similar to RECON [32] employs both contrastive and MAE [15] reconstruction losses with an additional multi-view image matching cost. These recent works typically rely on frozen text and image encoders, and all used point-cloud encoders. Our work shows that taking embeddings of multi-view images is more efficient and works better with real-world images.
Multi-view images for text-3D representations. Multi-view images have long been used to represent 3D shapes [2, 37, 40, 46, 13], with recent work [39, 24] using multi-view images for shape representation by aggregating CLIP features. While these works also use CLIP with multi-view images, our focus is whether using a trainable multi-view shape encoder as a drop-in-replacement for the point-cloud encoder in contrastive settings such as OpenShape will result in better shape-representation for downstream tasks. These works do not focus on improved representation learning, but on how to use prompting to select better views or view aggregation and improve shape classification. There is also a line of work [56, 61, 17] that uses multi-view depth maps from point cloud to adapt CLIP models for 3D. However, the focus of these methods is still on point clouds and the models show limited performance on large-scale datasets [50, 25].
3 Method
Figure 2: Our multi-view encoder takes a variable number $[1,M]$ of images as input and outputs a single shape embedding. The first few layers of the ViT model are frozen (blue) and attention operates on the individual frames in parallel. For the latter layers, the attention layers are trainable (orange) and modified to attend over all $M$ views. The embeddings for each frame are averaged to get the final embedding. Note that only major components of the model are depicted.
Our Duoduo CLIP learns a shape representation by using contrastive learning to encode multi-view images into a pre-aligned text-image space. Our contrastive learning framework is similar to previous works [49, 50, 25, 58, 60], with a shape encoder $E^{S}$ and frozen image $E^{I}$ and text $E^{T}$ encoders from a pre-trained CLIP model. The key difference of our model is that we replace the point cloud based shape encoder with one that uses multi-view images. Figure 1(a) shows a comparison of our model to prior work, and Figure 2 shows the details of our multi-view shape encoder.
To perform training, the dataset is organized into triplets of $(S_{i},I_{i},T_{i})$, where $S$ is the shape representation, $I$ are rendered images of the 3D object and $T$ is the corresponding text description. Embeddings are calculated for each modality $f^{S}_{i}=E^{S}(S_{i})$, $f^{I}_{i}=E^{I}(I_{i})$ and $f^{T}_{i}=E^{T}(T_{i})$. The asymmetric contrastive loss [57] between any two modalities is:
$l^{a\rightarrow b}_{i}=-\log\frac{\exp(\langle f^{a}_{i},f^{b}_{i}\rangle)/\tau}{\Sigma_{k=1}^{N}\exp(\langle f^{a}_{i},f^{b}_{k}\rangle)/\tau}$ (1)
where $a,b$ is any two modalities, $\tau$ is the temperature and $N$ is the batch size. Then, the full loss is:
$L_{CON}=\frac{1}{4N}\Sigma_{i=1}^{N}(l^{S\rightarrow T}_{i}+l^{T\rightarrow S}_{i}+l^{S\rightarrow I}_{i}+l^{I\rightarrow S}_{i})$ (2)
Multi-view shape encoder. For our multi-view shape encoder, we use rendered images of the object. Note that the input shape and image representations are the same, i.e. $S_{i}=I_{i}=\{r_{1},...,r_{m}\}$, where $r_{j}$ is a rendered frame, and $m$ is the number of multi-views. $E^{I}$ and $E^{S}$ share the same network architecture, with the main difference between $E^{I}$ and $E^{S}$ is that $E^{I}$ is frozen, while we train $E^{S}$. Weights for $E^{S}$ are initialized from $E^{I}$. Empirically, we found it was useful to keep the contrastive losses $l^{S\rightarrow I}_{i}+l^{I\rightarrow S}_{i}$. We suspect that by keeping $E^{I}$ frozen and the contrastive loss with the original image embeddings, our $E^{S}$ can learn a good representation that is close to the text embeddings, but does not drift too much away from the initial image embeddings.
For both $E^{S}$ and $E^{I}$ we average the extracted embeddings for all $m$ views to produce a single latent representation for an object: $f^{S}_{i}=E^{S}(S_{i})=\frac{1}{m}\Sigma_{j=1}^{m}\text{ViT}(S_{i})[j]$, where ViT [9] is the backbone encoder. Here we abuse the notation a bit, since our method allows interaction between views in the model, the index operation gets the corresponding embedding from the $j$-th frame’s [CLS] token. We initialize $E_{S}$ initially from $E^{I}$. During training, we randomly sample $1$ to $M$ views for each step so that the model is robust to a variable number of views.
Multi-view attention. We incorporate information flow between views by modifying the self-attention layers to attend over tokens across multiple views similar to recent multi-view aware methods like MVDream [38]. The original self-attention formulation [44] is $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}}V)$, where $Q,K,V\in R^{L\times k}$, and $L$ is the token length for a single image after patchify operations and $k$ is the channel dimension. For cross-view attention we aggregate token features across $m$ multi-views such that $Q^{\text{MV}},K^{\text{MV}},V^{\text{MV}}\in R^{(L\times m)\times k}$ (MV stands for multi-view) and the $\operatorname{Attention}(Q^{\text{MV}},K^{\text{MV}},V^{\text{MV}})$ can be calculated without changing the parameters of the original model. After attention, the views are separated and processed in parallel in other parts of the model. This mechanism helps the model reason between different views of the object and aggregate the most important information across all tokens.
Trainable layers. To reduce the memory needed for fine-tuning, we only train the self-attention layers. We further cut down on memory usage by reducing the number of trainable layers. Since the non-trainable self-attention layers have only been trained to see patches of a single image, we only enable the multi-view attention layers for the unfrozen attention layers (see Figure 2).
Figure 3: Left. We compare the Top 1 accuracies on Objaverse-LVIS for our model (Duoduo CLIP) and two pretrained CLIP models (zero shot) for different number of frames. B-32 is the model we initialize from and train with and BigG-14 is the model used by OpenShape. The dashed lines show the difference between using 1 and 12 frames. Right. We show the Top1, Top3 and Top5 accuracies for our model for 1 to 12 frames (note the y-axis starts at 40). Gray dashed lines shows the inference time(s) for 1 iteration and batch size 200 at different number of frames.
4 Experiments
To compare our Duoduo CLIP to prior work, we conduct experiments showing shape classification on a collection of 3D assets (on Objaverse-LVIS in Section 4.2), real-world multi-view images (on MVImgNet in Section 4.3), and more fine-grained text-to-shape retrieval (Section 4.4). We first describe the experimental setup including implementation details and the datasets we use.
4.1 Experimental setup
[…]
3D Datasets. We follow OpenShape [25] and train our model on a combination of four datasets consisting of Objaverse [7], ABO [6], ShapeNet [3] and 3D-FUTURE [12]. Text for the shapes consists of captions produced by image-captioning models and filtered text descriptions from metadata of the 3D objects crawled from the web (see Liu et al. [25] for details). In total, the combined dataset contains $\sim$874k shapes, with 46k shapes within the LVIS subset of Objaverse which are used for evaluation. As in OpenShape [25], we consider two variations of the data: 1) Ensembled which contains all shapes, and 2) Ensembled (no LVIS) which excludes shapes from the LVIS subset.
Renderings. We use images from Zero123 [26], which provides renderings of most shapes from Objaverse from $12$ random views using spherical sampling. The shapes not included in Zero123 (a few remaining objects in Objaverse and the other 3 datasets) are rendered using their Blender  [^1] https://www.blender.org/ script to match the shapes in OpenShape. We render $12$ views for all these objects.
Real-world data. MVImgNet [53] consists of multi-view images for 220k real-world objects over 238 categories. We first filter objects that has at least $12$ multi-views to obtain $\sim$190k remaining objects, where $12$ views are evenly sampled. BLIP-2 [23] is then used to produce captions for each of the views. To compare with point cloud based methods, we use the MVPNet subset containing $\sim$87k objects with densely reconstructed point clouds. After the preprocessing and filtering above, we obtain from MVPNet, $\sim$66k objects for the train split and $\sim$16k objects for the validation set, which has 180 categories. Note that the objects in the training and validation splits for MVPNet are disjoint but are both included in MVImgNet.
Table 2: Classification accuracy on Objaverse-LVIS trained on Ensembled (no LVIS) and Ensembled. The zero shot CLIP (B-32 and BigG-14) models are not trained on 3D data and serve as baselines. For each model, we indicate the representation used for computing the embeddings, multi-view (MV) or point-cloud (PC), and the shape encoder (Enc) architecture that is used. Our multi-view based Duoduo CLIP is able to outperform point-cloud based approaches with the exception of Uni3D is a large model with 1B parameters. $\dagger$ indicates our reimplementation of OpenShape which is faster and outperforms the original OpenShape.
Ensembled (no LVIS) Ensembled
Method Rep Enc Top1 Top3 Top5 Top1 Top3 Top5
ZS B-32 (12F) MV Avg 35.7 54.8 62.1 35.7 54.8 62.1
ZS BigG-14 (12F) MV Avg 43.5 64.2 71.3 43.5 64.2 71.3
ULIP [49] PC PointBERT 21.4 38.1 46.0 26.8 44.8 52.6
OpenShape [25] PC PointBERT 39.1 60.8 68.9 46.8 69.1 77.0
[…]
TAMM [58] PC PointBERT 42.0 63.6 71.7 50.7 73.2 80.6
Uni3D [60] PC 3D VIT 47.2 68.8 76.1 55.3 76.7 82.9
ShapeLLM [33] PC RECON++ 53.7 75.8 82.0
VIT-LENS [22] PC VIT-LENSG 50.1 71.3 78.1 52.0 73.3 79.9
Duoduo CLIP (5F) MV MVA 51.3 73.1 79.9 54.1 76.0 82.4
Duoduo CLIP (12F) MV MVA 52.7 74.5 81.3 55.2 77.1 83.4
4.2 Classification on Objaverse-LVIS
We conduct classification experiments on Objaverse-LVIS to investigate the performance of different variants of Duoduo CLIP and compare our model to prior work. As default settings for our experiments in this section, we use $12$ multi-views to encode an object for evaluation. And only the multi-view attention of the latter 6 layers of the model is trained (total 12 layers). We run evaluations with 3 different seeds and average the results for experiments. Unless otherwise stated models are trained with the Ensembled dataset. For classification, we use do nearest neighbor search of the multi-view shape embedding with class embeddings. As in OpenShape, for each class label, we construct several sentences with the label and then take the average of the sentence embeddings as the class embedding. The final class chosen is the class with the highest average similarity between the class embedding and the shape embedding.
We compare the performance of our model against two CLIP models for different numbers of frames in Figure 3, and recent state-of-the-art models in Table 2. For the CLIP models, we selected VIT with B-32 (the model that our Duoduo CLIP is initialized with), and BigG-14 (the CLIP teacher model used by OpenShape). Both models are run in a zero-shot manner without training on additional 3D data (the weights are frozen, with embeddings from the image encoder averaged).
How well does performance scale with number of views? Having a more complete view of a 3D object is important for 3D representation learning as elements of the object may be occluded for certain views. Thus we investigate how the number of views affects the classification accuracy. From Figure 3 (left), we see that our model performance scales better as the amount of information (views) increases compared to the zero-shot CLIP models ($11.22$ for our Duoduo CLIP, and $7.45$ for B-32, the model we used for initializing our weights). This suggests that our training strategy of randomly sampling 1 to 6 views during training is effective in letting the model learn to leverage the number of views given to it. Although, the accuracy increases as the number of multi-views increases, the inference time also increases (Figure 3 right). As the accuracy tapers off as we increase the number of views, to avoid additional computational cost during inference, we use 12 views as the default.
How many multi-views to beat a point cloud? Compared with using multi-views, using a point cloud representation offers good coverage of the entire geometry of the 3D object. Thus, we ask the question of how many views does it take to match existing methods on the Objaverse-LVIS classification task. We compare against ULIP [49], OpenShape [25], TAMM [58] that use point clouds with PointBERT [54] as the 3D encoder, but different ways of fine-tuning the 3D encoder. We also compare against more recent methods [60, 33, 22] that are the current top-performing methods on Objaverse-LVIS. From Table 2, on the Ensembled dataset, we are able to beat most previous methods with 5 multi-views except for Uni3D which has a much larger model (1B parameters). When we increase to 12 frames we are on par and even surpass Uni3D. This is also while using much fewer resources as seen in Table 1.
Does multi-views generalize better compared to point cloud? From the Ensembled (no LVIS) column of Table 2, we see that even the zero-shot BigG-14 model used as the teacher model for OpenShape beats it. Suggesting that the CLIP models already generalize well to unseen data, and point cloud methods potentially having worse generalization on unseen shapes. With our model that is further trained on Ensembled (no LVIS) we surpass point cloud methods by a large margin. Showing that fine-tuning the CLIP model better preserves their generalization ability.
Table 3: Multi-view attention (MVA) ablation of accuracy on Objaverse-LVIS.
Method Frames Top 1 Top 3 Top 5
-MVA 6 54.22 76.13 82.74
+MVA 6 54.59 76.46 82.89
-MVA 12 55.02 76.92 83.49
+MVA 12 55.32 77.08 83.49
Table 4: Number of layers ablation. Accuracy on Objaverse-LVIS.
Method Top 1 Top 3 Top 5
3 layers 53.77 75.8 82.41
6 layers 55.24 77.08 83.43
12 layers (full) 55.32 77.08 83.49
Does multi-view attention help? We verify the effectiveness of our cross view attention in Table 4. For both -MVA and +MVA, we train all 12 attention layers, with the difference being +MVA has the cross-views attention enabled. We see that using multi-view attention helps the performance most in boosting Top1 accuracy for both 6-frame and 12-frame evaluations.
Do we need to train all the layers? We conduct an ablation to see if it is necessary to train all 12 layers. In Table 4, we show the accuracy for restricting the training to the last 3 and 6 MVA layers vs the full 12 MVA layers. We see that the number of trainable layers can be cut by half with minimal losses to accuracy, while further cuts start to hurt model performance. Training only half the layers reduces the VRAM usage by roughly half and allows us to train on GPUs with less memory. By restricting the training to the last 6 MVA layers, we can shift from training on 4$\times$×A40s (48 GB) to 4$\times$×A5000s (24 GB) (see Table 1).
Table 5: MVImgNet classification comparison.
Method Top 1 Top 3 Top 5
ZeroShot B-32 (12F) 52.68 70.99 77.22
[…]
Ours (12F) 49.16 66.96 74.12
Ours (+MVPNet) (1F) 59.23 76.12 81.74
Ours (+MVPNet) (6F) 64.18 80.80 85.92
Ours (+MVPNet) (12F) 64.44 81.11 85.97
Ours (+MVImgNet) (1F) 61.75 79.07 84.18
Ours (+MVImgNet) (6F) 65.70 82.44 87.08
Ours (+MVImgNet) (12F) 66.06 82.72 87.21
Table 6: Text to shape retrieval comparison. Methods are trained on the Text2Shape [4] dataset (T2S) or Ensembled.
Model Data RR@1 RR@5
Text2Shape [4] T2S 0.40 2.37
Y2Seq2Seq [14] T2S 2.93 9.23
Parts2Words [41] T2S 12.72 32.98
TriCoLo [36] T2S 12.22 32.23
MXM-CLR [45] T2S 16.83 39.06
ZeroShot B-32 Images 11.90 27.85
[…]
Uni3D-G Ens. 10.78 26.37
Duoduo CLIP Ens. 15.90 34.08
4.3 Classification on real-world data
To investigate how well our Duoduo CLIP works on real-world images, we evaluate on multi-view images from MVImageNet. For these experiments, we take our Duoduo CLIP model trained on the Ensembled dataset and add additional training data from MVImgNet. We consider two sets of data from MVImgNet, the MVPNet training split of objects or the entire MVImgNet (after our preprocessing and filter). For comparison with point cloud based methods, we evaluate on the validation set of MVPNet. As a representative of point cloud based methods, we choose OpenShape as it is relatively resource-light compared to other point cloud methods.
Classification results. We report accuracy results on MVPNet in Table 6. For OpenShape, the top-1 accuracy is extremely low (only $3.55\%$ top-1) when evaluating on MVPNet directly with the model trained on Ensembled only. We suspect this is due to noisy reconstructed point clouds and inconsistent orientations compared to the training data. After training on MVPNet data, OpenShape’s performance improves significantly. For our Duoduo CLIP, even without training on MVImgNet or MVPNet data, our model achieves a top-1 accuracy of $49.16\%$. However, this is lower than the zero-shot CLIP B-32 model that our model is initialized from, suggesting possible overfitting to synthetic data. When we include training with multi-view images from MVPNet, our method surpasses both OpenShape and the zero-shot CLIP B-32 model even when just using one 1 view. Performance is further boosted when trained on the entire MVImgNet, highlighting the advantages of our Duoduo CLIP which uses multi-view images over OpenShape’s point clouds. By using multi-view images, we have better generalization on unseen data due to the priors from initializing with CLIP B-32. We also can train on more data, as multi-view images are more readily available compared to point clouds.
Figure 4: Retrieval results for OpenShape and Duoduo CLIP given input images from MVPNet. We perform retrieval with 12 input images, on the left most column we show one of the frames of the query object. For both methods we show the top 3 retrieval results using Objaverse as a retrieval library. Overall, Duoduo CLIP retrieves shapes that are more similar to the query image, while OpenShape sometimes retrieves broken geometry (row 3).
Retrieval examples. In Figure 4, we present example retrieval results for querying Objaverse objects using images from the MVPNet validation set. We compare retrievals using OpenShape (+MVPNet) and Duoduo CLIP (+MVPNet). For OpenShape, we use 12 query images for each object, encode them with the frozen BigG-14 CLIP model, and average the image embeddings to obtain the query embedding. [^2] For the largest BigG-14 CLIP model, OpenShape does not use any linear projection layer $g^{I}$. This embedding is then matched against a database of point cloud embeddings of Objaverse objects to retrieve the top 3 results. For our method, Duoduo CLIP encodes both the query and Objaverse objects using 12 multi-view images.
From Figure 4, we see that for basic objects like benches, both methods are able to retrieve plausible objects. For the armchair (row 2), while OpenShape retrieved objects of the same semantic class, our Duoduo CLIP retrieved armchairs that has better matching textures. This is because multi-view images can represent textures better than point clouds. Note that some retrieved objects look similar due to duplicate objects in Objaverse. For the jacket (row 3), the objects returned by OpenShape have broken geometry and are incoherent. We hypothesize this may happen when the sampled points produce embeddings that are by chance similar to the query image. This issue is also present last example (a red and black soccer ball) where the first object retrieved by OpenShape has broken geometry. However, our method also fails in this case, as our model retrieves ladybugs with red wings and black dots that resemble the soccer ball. Interestingly, OpenShape is able to retrieve a soccer ball and a red ball. This can indicate that point cloud based methods may be more robust to geometric properties of the object compared to the rendered images. However, we find that overall the retrieved objects from our method is more consistent.
4.4 Fine grained text-to-shape retrieval
Previous evaluations focused on classification or image-based retrieval and do not assess how well the methods align fine-grained text descriptions with 3D models. To evaluate this aspect, we use the test split of Text2Shape [4], which contains $\sim$1.4k tables and chairs from ShapeNet and $\sim$7.4k human-captioned descriptions ($\sim$5 per shape). We report the shape retrieval rates at 1 and 5.
Retrieval performance. We compare the performance of our model to prior work in Table 6. We also include previous state-of-the-art methods that were trained on the Text2Shape dataset. Overall, compared to other methods trained on the Ensembled dataset, our method demonstrates superior performance. The zero-shot performance alone is impressive, surpassing point cloud methods. Notably, we observed a significant discrepancy between the zero-shot CLIP model reported by Ruan et al. [36] and our findings, where the LAION checkpoint trained by OpenCLIP significantly outperforms the original OpenAI checkpoint. Lastly, our model approaches the performance of the SOTA method trained directly on Text2Shape, indicating that there is still much to explore regarding more fine-grained text alignment in large-scale text-to-3D models.
Figure 5: Retrieval using fine-grained text from the Text2Shape dataset. On the left is the input text description and we show the top 3 retrieval results for Uni3D and Duoduo CLIP with the ground truth shape highlighted in green.
Retrieval examples. We present example retrieval results compared to Uni3D in Figure 5. Generally, for text descriptions emphasizing complex textures, such as the ‘man image’ in the first row, our model successfully retrieves the corresponding ground truth 3D model. However, for descriptions focusing on geometric properties, like ‘round back’ in the second row, point cloud-based methods tend to retrieve shapes that better match the specified criteria.
5 Limitations and Future Work
We introduced Duoduo CLIP, an efficient joint text and 3D representation method that leverages multi-view images and surpasses current point cloud based state-of-the-art methods across almost all quantitative metrics while using significantly fewer resources on both synthetic and real datasets.
However, our method sometimes fails in cases requiring geometric understanding, as seen in the qualitative retrieval results from images or text. This is an inherent advantage of point clouds, as each point contains three-dimensional coordinates. To address this, we could explore additional inputs to the model such as depth or normal images of the objects, which provide more geometric information, especially with recent depth estimation methods becoming better. Another potential way to improve performance is with LLMs, as demonstrated by Tang et al. [42], where representations from ULIP-2 were used by an LLM to boost performance and enable question-answering applications. Using our model might further increase performance, offering the flexibility of requiring only a single or a sparse number of images as input. Banani et al. [1] also point out the relative weakness of CLIP pretraining on both 3D structure and consistency benchmarks compared to newer SSL methods [59, 30], which can potentially replace the current pretraining framework used by most works.
Overall, we believe that our method will provide better resource scaling compared to point cloud based methods as more 3D data becomes available in the future.
Acknowledgements. This work was funded by a CIFAR AI Chair, a NSERC Discovery grant, and a CFI/BCKDF JELF grant.